PRE-LAB EXERCISES average size. If d is the diameter in millimeters, some pebbles will be bigger and some smaller than the desired average do. Suppose you have a rock crusher. The pebbles it makes have different sizes, but you can adjust the 1. (a) In the histogram below, about how many pebbles are shown to have diameter of greater 150 100 (b) In this distribution (same histogram,) approximately how many pebbles are represented 50 altogether, including all sizes? 20 40 60 80 Pebble Diameter d in Millimeters

e you made a batch of 7 x107pebbles crushed by the same rock crusher, using the same same raw material. Estimate how many pebbles would have diameter of greater (c) Suppos settings and the Remember to explain your reasoning in sentences.) (d) Let do represent the average pebble diameter. About half of the pebbles will have diameters less than do and about half will have greater diameter, but this is not always precisely true. For example, consider the set of numbers {1,2,3,4,5,15). Calculate the mean or average value and show that this value does not fall in the middle of the distribution, where middle means the middle number or median. About where would the middle be?

Nevertheless, it is often reasonable to suppose the mean value is above approximately at the center of the set of values. (e) Let us consider the spread of distribution in pebble size. How much spread is there around the mean value? You cant just average the difference d -do for each data point, because that average is always zero. So, what you do first is square the differences, then average the s (all positive numbers) and then take the square root. Here we write this quantity A, size N is large enough usually the symbol ơ the standard deviation. Actually, A is a root mean square difference. The quantity befo the square root is called variance. So, loosely speaking Δ quares if the sample is used because it is approximately what they call is the standard deviation. For o ur 2-2 UND Physics CPSL

Nature of Measurement Uncertainty purposes, lets just call A the rms for short. It turns out that when the distribution looks sort of bell-shaped, about 68% of the data fall between d- and d, +A, Also about 95% of the data fall between d, -2A and d +2A. So, the point of the rms (or in other words A) is that it measures how spread out the distribution is. For the following (short!) list of pebble sizes, find the mean value and the rms (Sizes d in mm) 1 5.846 2 10.057 3 4.98 4. 7.104 5. 7.856 6. 9.392 7 5.927 8 12.301 9 10.159 10 9.507 11 3.442 12. 3.476

(f) Compose two sentences to give the practical meaning (more than a definition) of the mean value and the rms of some data. In other words, what does each of these two statistics actually tell you about the data? The next problem also has to do with multiple measurements. It has to do with time measurements rather than diameter measurements, but the ideas are the same. From your reading, write down formulas for the distance traveled and for the velocity, each as a function of time, in terms of initial speed and the acceleration a, assuming the acceleration is constant. This situation applies to a falling object or a car skidding to a stop 2.

Need help with 1: b, c, d, e and f. And 2. Thanks!
Answer #1

1.

(b)

Number of pebbles represented by the given histogram (approximately) are: 10+20+40+90+110+140+145+140+110+100+50+35+15+10 =1,015​​​​​​ pebbles.

(c)

Given: X =65 mm.

Frequency table (including the relative frequency and cumulative relative frequency) of the above histogram is:

1 Class Frequency(f) Relative frequency Cumulative relative frequency 220-25 з 25-30 4 30-35 5 35-40 640-45 7 45-50 850-55 955-60 10 60-65 11 65-70 12 70-75 13 75-80 14 80-85 15 85-90 16 TOTAL1015 17 10 0.009852217 20 0.019704433 40 0.039408867 90 0.088669951 110 0.108374384 140 0.137931034 145 0.142857143 140 0.137931034 110 0.108374384 100 0.098522167 50 0.049261084 35 0.034482759 15 0.014778325 0.009852217 0.02955665 0.068965517 0.157635468 0.266009852 0.403940887 0.54679803 0.684729064 0.793103448 0.891625616 0.9408867 0.975369458 0.990147783 + 10 0.009852217

Sigma f =1015; Relative frequency =f/Sigma f; Cumulative relative frequency = Total relative frequency upto that class (leq upper limit of the respective class).

From the 'Cumulative relative frequency' column of the above table (shown in Yellow shading):

P(Xleq65 mm) =0.793103448 (here P is the Probability).

Now, P(X>65 mm) =1 - P(Xleq65 mm) = 1 - 0.793103448 =0.206896552

Hence, the number of pebbles out of 7x107 pebbles that would have diameter of grater than 65 mm =

7x107​​x0.206896552 =14482759 (rounded off).

(d)

Mean of given set of numbers =(1+2+3+4+5+15)/6 =30/6 =5

The mean value of 5 is not in the middle because middle values are 3 and 4, that is, median =(3+4)/2 =3.5

(e)

​​​​Sample mean of the given data, ar{X} =ΣΧ/n 90.047/12 =7.504 mm

Sample Std.deviation, s =Σ(X-X)2/(n-1)2.841 mm

(Here, we use '(n-1)' instead of 'n' because it removes any bias).

(f)

On an average, the size of a pebble is 7.504 mm.

On an average, the deviation of a pebble from the average value of 7.504 mm is 2.841 mm. (that is, the average deviation of a pebble from its mean size is 2.841 mm and it may be less than or greater than the mean).